# Systems with Switching Causal Relations: A Meta-Causal Perspective

### v1

@@ 5 Related Work
Causal Transportability and Reinforcement Learning. Meta-causal models cover cases that reduce to conditionally dependent causal graphs due to changing environments (Peters et al., 2016; Heinze-Deml et al., 2018), but also extend beyond that for dynamical systems. In this sense, the work of Talon et al. (2024) takes a meta-causal view by transporting edges of different causal effects between environments. In general, the transportability of causal relations (Zhang & Bareinboim, 2017; Lee & Bareinboim, 2018; Correa et al., 2022) can be thought of as learning identification functions that have identified the general conditions of the underlying processes to transfer certain types of causal effects between environments Sonar et al. (2021); Yu et al. (2024); Brouillard et al. (2022). This has been studied to some extent under the name of meta-reinforcement learning, which attempts to predict causal relations from the observations of environments Sauter et al. (2023); Dasgupta et al. (2019). Generally, transferability has been considered in reinforcement learning, where the efficient use of data is omnipresent and causality provides guarantees regarding the transferability of the observed mechanisms under changing environments (Sæmundsson et al., 2018; Dasgupta et al., 2019; Zeng et al., 2023).
@@
Large Language Models (LLMs). Meta-causal representations are an important consideration for LLMs and other foundation models, since these models are typically limited to learning world dynamics from purely observational textual descriptions. In order to reason in a truly causal manner, LLMs need to learn meta-causal models (sometimes referred to as ‘world models’) that allow them to simulate the implied consequences of interventions (Lampinen et al., 2024; Li et al., 2021; 2022). To the best of our knowledge, Zečević et al. (2023) made the first attempt to define explicit meta-causal models that integrate with the Pearlian causal framework. However, their proposed models are defined purely as adjacency matrix memorization, such that causal reasoning in LLMs is equivalent to knowledge recall: $(X\rightarrow Y)\Longleftrightarrow(e_{XY}\in\text{Mem})$. Thus, LLMs would not learn the underlying mediation process and would fail to predict the effects of novel interventions on the causal graph.
@@
Actual Causality, Attribution and Cognition. Our MCM framework can also be used to infer and attribute responsibility (as shown in Sec. 4.1) and therefore touches on the topics of fairness, actual causation (AC) and work on counterfactual reasoning in cognitive science. (Von Kügelgen et al., 2022; Karimi et al., 2021; Halpern, 2000; 2016; Chockler & Halpern, 2004; Gerstenberg et al., 2014; Gerstenberg, 2024). The similarities extend to how MCM encourage reasoning about the dependence between actual environmental contingencies and qualitative types of causal mechanisms. In this sense, MCM allow for the direct characterization of actual causes of system dynamics types. A rigorous formalization is left open for future work. While AC in combination with classical SCM only describes relationships between causal variable configurations and observable events, there are cases where we can take an MCM and derive an SCM that encodes the meta-causal types with instrumental variables, thus allowing a similar meta-causal analysis within the AC framework.
@@ 6 Conclusion
We formally introduced meta-causal models that are able to capture the dynamics of switching causal types of causal graphs and, in many cases, can better express the qualitative behavior of a causal system under consideration. Within MCM typing mechanisms generalize the notion of specific structural equations and abstract away unnecessary details. We presented a first, motivating example of how a classical causal analysis and a meta-causal perspective can attribute different root causes to the same situation. We extended these claims by considering that MCM are still able to capture and attribute changes in the mechanistic behavior of the system, even though no actual change is apparent. As an application, we presented a first approach to recover metacausal states in the bivariate case. Although our experimental results represent only a first preliminary approach, we find that meta-causal models are a powerful tool for modeling, reasoning, and inferring system dynamics. Finally, we have shown how MCM are able to express more information about a dynamical system than a simple conditioning variable in a standard SCM can do.
Limitations and Future Work. While we have formally introduced meta-causal models, there remain several practical open directions to pursue. First, we have been able to provide examples that illustrate the differences between standard causal, and meta-causal attribution. In particular, the joint application of standard and meta-causal explainability will allow for the joint consideration of actual and mechanistic in future attribution methods. Second, our approach to recover meta-causal states from unlabeled data is open to extension. Discovery on the full causal graphs is a desirable goal that, however, needs further work to reliably control for the changing sets of parents under changing MCS. Finally, we have made a first attempt to present examples for and against the controlling or observational role of meta-causal models. While the observational perspective is mainly of interest for analytical applications, the application of meta-causal models to agent systems and reinforcement learning opens up a new field of application.
@@ Acknowledgments
The TU Darmstadt authors acknowledge the support of the German Science Foundation (DFG) project “Causality, Argumentation, and Machine Learning” (CAML2, KE 1686/3-2) of the SPP 1999 “Robust Argumentation Machines” (RATIO). This work is supported by the Hessian Ministry of Higher Education, Research, Science and the Arts (HMWK; projects “The Third Wave of AI”). It was funded by the European Union. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. Grant Agreement no. 101120763 - TANGO. This work was supported from the National High-Performance Computing project for Computational Engineering Sciences (NHR4CES).
@@
In Sec. A we derive the formula for the theoretical upper bound of the $95\%$ confidence interval. Sec B discusses the sample statistics and convergence results of the applied LO-RANSAC algorithm. Finally, Sec. C discusses the exact parameters of the bivariate meta-causal mechanism data sampling.
Appendix A Probabilities for Sample Computation and Upper Bounds
@@
Appendix B EM Convergence Results
The required number of resamples for a $95\%$ success rate of the RANSAC algorithm is calculated by $\text{log}(0.05)/\text{log}(1-C^{1})$, where $S1$ are the convergence rates for the individual samples computed in Sec. A.
@@ Appendix B EM Convergence Results
Appendix C Mechanism Sampling
@@
4:         $\text{N}\leftarrow\text{requiredSamples}(k,\text{maxClassDev},0.95)$triangleright Compute ##\## of samples (c.f. Sec. A)

### v2

@@ 5 Related Work
Causal Transportability and Reinforcement Learning. Meta-causal models cover cases that reduce to conditionally dependent causal graphs due to changing environments (Peters et al., 2016; Heinze-Deml et al., 2018), but also extend beyond that for dynamical systems. In this sense, the work of Talon et al. (2024) takes a meta-causal view by transporting edges of different causal effects between environments. In general, the transportability of causal relations (Zhang & Bareinboim, 2017; Lee & Bareinboim, 2018; Correa et al., 2022) can be thought of as learning identification functions that identify general conditions of the underlying processes to transfer certain types of causal effects between environments Sonar et al. (2021); Yu et al. (2024); Brouillard et al. (2022). This has been studied to some extent under the name of meta-reinforcement learning, which attempts to predict causal relations from the observations of environments Sauter et al. (2023); Dasgupta et al. (2019). Generally, transferability has been considered in reinforcement learning, where the efficient use of data is omnipresent and causality provides guarantees regarding the transferability of mechanisms between environments (Sæmundsson et al., 2018; Dasgupta et al., 2019; Zeng et al., 2023).
Gating Models. Modeling switching causal relations (Liu et al., 2023) via so called ‘gates’ has been considered in prior works Minka & Winn (2008); Winn (2012). While MCMs extend beyond context-specific independencies, gates pose a practical way of modeling switching relations in cases where MCMs can be reduced to context-conditioned SCM.
@@
Large Language Models (LLMs). Meta-causal representations are an important consideration for LLMs and other foundation models, since these models are typically limited to learning world dynamics from purely observational textual descriptions. LLMs need to learn meta-causal models that allow them to simulate the consequences of interventions (Lampinen et al., 2024; Li et al., 2021; 2022). To the best of our knowledge, Zečević et al. (2023) made the first attempt to define explicit meta-causal models that integrate with the Pearlian causal framework. However, their MCM are purely defined as adjacency matrix memorization, such that causal reasoning in LLMs equals a simple knowledge recall $(X\rightarrow Y)\Leftrightarrow(e_{XY}\in\text{Mem})$ and thus fails to generalize to novel scenarios.
@@
Actual Causality, Attribution and Cognition. Our MCM framework can be used to infer and attribute responsibility, as shown in Sec. 4.1, and therefore touches on the topics of fairness, actual causation (AC) and work on counterfactual reasoning in cognitive science. (Von Kügelgen et al., 2022; Karimi et al., 2021; Halpern, 2000; 2016; Chockler & Halpern, 2004; Gerstenberg et al., 2014; Gerstenberg, 2024). The similarities also extend to how MCM encourage reasoning about the dependence between actual environmental contingencies and qualitative types of causal mechanisms. In this sense, MCM allow for the direct characterization of actual causes of system dynamics types, but a rigorous formalization is open for future work. While AC in combination with classical SCM only describes relationships between causal variable configurations and observable events, there are cases where we can take an MCM and derive an SCM that encodes the meta-causal types with instrumental variables, thus allowing a similar meta-causal analysis within the AC framework.
@@ 6 Conclusion
We formally introduced meta-causal models that are able to capture the dynamics of switching causal types of causal graphs and, in many cases, better express the qualitative behavior of the system under consideration. Within MCM, types generalize the notion of specific structural equations and abstract away unnecessary detail. We presented a motivating example of how a classical causal and a meta-causal inference might disagree on the attribute of root causes. We extended claims by considering that MCM are still able attribute changes in mechanistic behavior of a system, even when no actual changes becomes apparent. We presented a first approach to recover meta-causal states in the bivariate case. Although our experimental results only represent a first preliminary approach, we find that MCM are a powerful tool for modeling, reasoning, and inferring system dynamics. We demonstrated how MCM can be deployed dynamical systems and proved that they extend beyond conventional SCM.
Limitations and Future Work. While we have formally introduced meta-causal models, there remain several open directions to pursue, which we briefly touch on in Appendix H. We have been able to provide examples that illustrate the differences between standard causal, and meta-causal attribution. In particular, the combined application of standard and meta-causal explainability will allow for the joint consideration of actual and mechanistic in future attribution methods. However, our approach to recovering meta-causal states from unlabeled data is open to extension. Discovery on the full causal graphs is a desirable goal that is difficult to achieve for the reasons discussed in this paper. Finally, we made a first attempt to present examples for and against the controlling or observational role of meta-causal models, which we briefly discuss further in Appendix I. While, the presented observational perspective on MCM is mainly of interest for analytical applications, the application to agent systems and reinforcement learning might open up further fields of application.
@@ Acknowledgments
The Technical University of Darmstadt authors received funding by the EU project EXPLAIN, funded by the German Federal Ministry of Education and Research (BMBF) (grant 01—S22030D). They acknowledge the support of the German Science Foundation (DFG) project “Causality, Argumentation, and Machine Learning” (CAML2, KE 1686/3-2) of the SPP 1999 “Robust Argumentation Machines” (RATIO). This work is supported by the Hessian Ministry of Higher Education, Research, Science and the Arts (HMWK; projects “The Third Wave of AI”). It was funded by the European Union. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. Grant Agreement no. 101120763 - TANGO. This work was supported from the National High-Performance Computing project for Computational Engineering Sciences (NHR4CES).
@@
The appendix is structured as follows: In Sec. A we describe the general emergence of causal relations in SCM from the underlying mediating process. In Sec. B we discuss desirable classes of identification functions. In Sec. C we consider the existence and compatibility of decomposed equations. In Sec. D we prove a condition for the reduction for meta-causal models to conditioned SCM. In Sec. E we derive the formula for the theoretical upper bound of the $95\%$ confidence interval. Sec F present the sample statistics and convergence results of the applied LO-RANSAC algorithm. In Sec. G we present the exact parameters for sampling bivariate meta-causal mechanism data. In Sec. H we discuss further practical applications of meta-causal models. Finally, in Sec. I, we discuss possible assertive properties of meta-causal models.
Appendix A Emergence of Causal Effects from Mediating Processes
The definition of a meta-causal frames (Def. 1) grounds the emergence of standard SCM to an underlying mediating process $\operatorname{\mathcal{E}}$. In particular, for any meta-causal frame, particular causal interactions are read via the identification function. While the underlying mediation process is time-dependent, the resulting causal graph is the projection of all causal interactions within the process onto a graphical structure. Note, however, that the resulting SCM still preserves the sequence of causal interactions through the DAG induced partial ordering of variables. This ‘logical’ time ordering (induced via the partial order) can be seen as an abstraction of the underlying process.
Appendix B Desirable Classes of Identification Functions
Particular choices of different identification functions will result in different identified meta-causal states. However, from a classical causal perspective it might be desirable to choose particular classes of functions to identify faithful SCM. In particular, one could ask the question whether indirect effects are of interest and should also be identified by the identification function. Consider a scenario where three variables $X,Y,Z$ are considered as part of an to-be-identified SCM from an underlying process. In our scenario we identify the direct causal effects $X\rightarrow Y$ and $Y\rightarrow Z$ and –under the assumption that no further direct edges can be identified–, assume the graph to be faithful and we do not identify the indirect causal relation $X\rightarrow Z$. This however changes, once variable $Y$ is dropped out of the variable set $\operatorname{\mathbf{X}}$ of the SCM. Now, with the same underlying mediating process we want our identification function to identify $X\rightarrow Z$, which was omitted before. Generally, we assume that in most cases identification function that identify direct causal effects with regard to $\operatorname{\mathbf{X}}$ are the most common.
Following this rather high-level discussion, we frame the previous procedure in terms of the identification function and functional relation. Given all functional relations $\mathcal{X}^{S}$, the type decoder can determine whether the whole functional relation $\mathcal{X}_{j}$ between variables $X_{j}$ is mediated by some other (set of) intermediate mechanism(s) $\mathcal{X}_{\mathbf{z}}$ [^3] Note in this context, that we assume structural equations to be uniquely identifiable. This might be implemented via type-theoretic considerations in which every variable of the underlying process gets assigned a unique type, such that for every $s_{i}\in\operatorname{\mathcal{S}}_{i}$, $\forall i,j.i\neq j\operatorname{\Rightarrow}\operatorname{\mathcal{S}}_{i}\neq\operatorname{\mathcal{S}}_{j}$ hold. Even though two functions might be isomorphic they can be thus be differentiated via their domain and codomain. In practice, one might additionally consider the computations graph of the mediation process to uniquely differentiate between isomorphic functions. In situations where $X_{i}\rightarrow X_{\mathbf{z}}\rightarrow X_{j}$ the type encoder needs to identify whether the relation $X_{i}\rightarrow X_{j}$ is purely mediated via $\mathcal{X}^{S}_{\mathbf{z}}$ (meaning that all computational paths from $X_{i}$ to $X_{j}$ in $\mathcal{X}^{S}_{j}$ that does not passes through $\mathcal{X}^{S}_{\mathbf{z}}$) or some additional direct effect exists that are not ‘shielded’ by $\mathcal{X}^{S}_{\mathbf{z}}$. Conversely, whenever some $X_{k}\in X_{\mathbf{z}}$ of the intermediate set is dropped/marginalized from the causal variables, the mechanism $\mathcal{X}^{S}_{k}\subset\mathcal{X}^{S}_{j}$ is no longer identified as the mechanism of a causal variable within the overall $\mathcal{X}^{S}_{j}$ and (given that the remaining $\mathcal{X}^{S}_{\mathbf{z}\setminus\{k\}}$ does not shield $X_{i}$ from $X_{j}$) a direct arrow $X_{i}\rightarrow X_{j}$ can be inferred.
Given our meta-causal framework we, however, have the freedom to identify the indirect relations. In that case one might assign these edges a particular ‘[1-hop] indirection’ type, indicating it to be the result of an indirect relation via an (possibly unobserved) intermediate variable. Consider, how this notion of edge types over unobserved variables is already in use in CPDAGs –as for example produced by the PC or GES algorithms (Spirtes et al., 2000; Chickering, 2002)– to indicate undirected edges for which the orientation is currently unknown.
Appendix C Compatibility of Decomposed Equations
In this section, we provide a brief discussion on the existence and compatibility of functions $X$ as considered in Equation 1.
Since the conditioned SCM in Eq. 1 is modeled after an the particular causal model that exists under a meta-causal state indexed by $z$, it follows that a particular composition of functions $f^{z}_{ij}$ has to exist (and that the functions are compatible), since the full function $f^{z}_{j}$ exists for each meta-causal state. In a naïve approach, the order of composition $i\in\{1..N\}$ enforces a particular evaluation order of the functions, and in particular requires this order to be the same for every meta-causal state.
Generally, the presented function composition might not work, in the case that the individual functions get chosen badly from the start. A simple solution to overcome such problem is to assume compositions of lifted functions and assume their signatures to be compatible (which is always permitted due to the known existence of the composed $f^{z}_{j}$). Note that functions $f^{z}_{ij},f^{z^{\prime}}_{ij}$, whose functional type $t_{ij}$ did not change under a change of $z$ to $z^{\prime}$, to make signatures compatible.
Common handling of compositionality in SCM: The adjustment of signatures is in fact often considered in the case of compositional SCM, e.g. additive noise models, where the signature of an outer ‘merging function’ $f_{j}(f_{1j},\dots,f_{Nj})$, e.g. $\sum_{i\in pa(j)}f_{ij}$, is in fact adjusted based on the number of parents (or, otherwise, zero weight edges are incorrectly excluded from the parent set).
Appendix D Reduction of Meta-Causal Models to Conditioned SCM
In this section, we provide a condition under which meta-causal models can be reduced to ordinary SCM with structural equations conditioned on some external variable as related to Eq. 1.
Definition 4 (MCM Reducability).
For a given mediation process $\operatorname{\mathcal{E}}=(\operatorname{\mathcal{S}},\operatorname{\sigma})$ and abstraction $\varphi:\operatorname{\mathcal{S}}\rightarrow\operatorname{\bm{\mathcal{X}}}$ we call a meta-causal model $\operatorname{\mathcal{A}}=(\operatorname{\mathcal{T}}^{N\times N},\operatorname{\mathcal{S}},\delta)$ reducible to a conditioned SCM if there is some SCM over $\operatorname{\mathbf{X}}\cup\{Z\}$ where the types of functional dependencies between $X_{i},X_{j}\in\operatorname{\mathbf{X}}$ can be fully determined by $Z$.
Theorem D.1 (Specific Criterion for Meta-Causal Reducability).
If for a given mediation process $\operatorname{\mathcal{E}}=(\operatorname{\mathcal{S}},\operatorname{\sigma})$ and abstraction $\varphi:\operatorname{\mathcal{S}}\rightarrow\operatorname{\bm{\mathcal{X}}}$, a meta-causal model $\operatorname{\mathcal{A}}=(\operatorname{\mathcal{T}}^{N\times N},\operatorname{\mathcal{S}},\delta)$ all its transitions are loops, then it is reducible to a conditioned SCM.
Proof.
If a meta-causal model only has loops as transitions, the meta-causal types are independent of the modeled mediation process. While types can differ for different starting conditions in the environment, its transition process never results in a type change. Hence, meta-causal types induce an equivalence relation on $\operatorname{\mathcal{S}}$, compatible with $\sigma$ and we can introduce an exogenous variable $Z$ with $\mathcal{Z}=\operatorname{\mathcal{T}}^{N\times N}$ or, alternatively, with values $\mathcal{Z}$ for each connected component in $\operatorname{\mathcal{E}}=(\operatorname{\mathcal{S}},\operatorname{\sigma})$. ∎
We also see some potential to weaken this criterion and, therefore, find a more general condition for reducibility by further examination of the abstraction that links the mediating process with the causal variables for future work. The primary obstacle in this regard is that meta-causal models abstract away some information about structural equations, such that interventions on $Z$ might lead to a mismatch between the resulting SCM and the underlying mediation process.
Appendix E Probabilities for Sample Computation and Upper Bounds
@@
Appendix F EM Convergence Results
The required number of resamples for a $95\%$ success rate of the RANSAC algorithm is calculated by $\text{log}(0.05)/\text{log}(1-C^{1})$, where $S1$ are the convergence rates for the individual samples computed in Sec. E.
@@ Appendix F EM Convergence Results
Appendix G Mechanism Sampling
@@
Appendix H Practical Applications of Meta-Causal Models
As the main focus of our initial work on meta-causal model lies on providing a first, formal, definition of meta-causal models, we tried to approach MCM from a spectrum of different theoretical perspectives. In this Section, we provide a brief discussion on two possible practical fields of applications for MCM.
Health and Medicine. First, we would like to expand on our stress-induced fatigue example as it can not be reduced to a standard SCM, and provide an actionable perspective on MCM. While we still assume the underlying neuropsychological process to be more complex, with multiple interplaying factors to influence each other, we consider the same simplified model as presented in the paper. We now assume that some drug exists that is able to influence certain health related processes within the patient, such that the underlying –previously self-reinforcing– stress relation is unconditionally changed to a suppressing one. (Upon closer consideration, the previous intervention might constitute a meta-causal do-operator, as we detach the functional type from the underlying dynamics and fix its particular functional type.)
This perspective not only allows the forecast of system changes, but also yields an actionable model which can be actively steered between meta-causal states. To permanently treat a patient, one could consider the objective to reach a self-stabilizing meta-causal state. Note how this meta-causal objective might be different to that of a classical causal one, where stress levels would similarly be reduced, but no attention is placed on the (possibly unchanged) system dynamics, such that stress levels might rise up again after the intervention ends.
Economics. Second, recall that our MCMs are defined as finite state machines. Figuring out exact transition conditions that induce meta-causal state transitions also yield important insights on the volatility/stability of systems in terms of risk analysis and policy making. Such scenarios might commonly arise in economics, where relations in markets can change due to the sudden appearance of disrupting factors (e.g. a new competitor entering the market or a financial crisis) while effects might persist even with the disrupting factor having vanished.
Appendix I Assertive Meta-Causal Models
In this section we will briefly touch upon possible assertive properties of meta-causal models. In this initial paper we chose to utilize a very general definition of types, which intentionally held flexible to allow for the most descriptive models. As for this definition, there might exist a gap between between the descriptive modeling of MCM and their ‘assertive’, data generating, properties. This gap primarily stems from the mapping of specific structural equations onto abstract types, which prevents a back projection of types to structural equations in the general case. With the special class of conditionally reducible SCM we, however, presented a particular class of MCM that yield ‘assertive’ properties by being able to translate types back onto the level of structural equations. Further restrictions on types and their relations to structural equations might be placed in specific applications, to shield users from misuse of the framework. Still, we where able to present several relevant applications of MCM even with this this most general formalization of MCM.
While meta-causal states might be mapped back to configurations of ‘classical’ causal models, we highlight that MCM are primarily concerned with the modeling of the overarching meta-causal state transitions. By modeling MCM as (possibly non-deterministic) finite-state machines, we, in fact, can make predictions about a systems’ future course on the meta-level. This includes the sampling of state trajectories and making assertions about their stability and similar properties.
@@ Appendix I Assertive Meta-Causal Models
4:         $\text{N}\leftarrow\text{requiredSamples}(k,\text{maxClassDev},0.95)$triangleright Compute ##\## of samples (c.f. Sec. E)